# Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities

### v2

@@ Abstract.
Security vulnerabilities in modern software are prevalent and harmful. While automated vulnerability detection tools have made promising progress, their scalability and applicability remain challenging. Recently, Large Language Models (LLMs), such as GPT-4 and CodeLlama, have demonstrated remarkable performance on code-related tasks. However, it is unknown whether such LLMs can do complex reasoning over code. In this work, we explore whether pre-trained LLMs can detect security vulnerabilities and address the limitations of existing tools. We evaluate the effectiveness of pre-trained LLMs, in terms of performance, explainability, and robustness, on a set of five diverse security benchmarks spanning two languages, Java and C/C++, and covering both synthetic and real-world projects.
@@
Overall, all LLMs show modest effectiveness in end-to-end reasoning about vulnerabilities, obtaining an average of 60% accuracy across all datasets. However, we observe that LLMs show promising abilities at performing parts of the analysis correctly, such as identifying vulnerability-related specifications (e.g., sources and sinks) and leveraging natural language information to understand code behavior (e.g., to check if code is sanitized). Further, LLMs are relatively much better at detecting simpler vulnerabilities that typically only need local reasoning (e.g., Integer Overflows and NULL pointer dereference). We find that advanced prompting strategies that involve step-by-step analysis significantly improve performance of LLMs on real-world datasets (improving F1 score by up to 0.25 on average). Finally, we share our insights and recommendations for future work on leveraging LLMs for vulnerability detection.
1. Introduction
@@
Security vulnerabilities afflict software despite decades of advances in programming languages, program analysis tools, and software engineering practices. Even well-tested and critical software such as OpenSSL, a widely used library for applications that provide secure communications, contains trivial buffer overflow vulnerabilities, e.g., (CVE-2022-3602, 2022) and (CVE-2022-3786, 2022). A recent study by Microsoft showed that more than 70% of the vulnerabilities are still caused by well-understood memory safety issues (Miller, 2019). The size and complexity of modern software systems are growing quickly, encompassing numerous programs, libraries, and modules that interact with each other. Hence, we need major technical advances to effectively detect security vulnerabilities in such complex software.
@@
Traditional techniques for automated vulnerability detection, such as fuzzers (Manes et al., 2018), and static analyzers such as CodeQL (Avgustinov et al., 2016) and Semgrep (Semgrep, 2023) have made promising strides. For example, in the last two years, researchers found over 300 security vulnerabilities through custom CodeQL queries (Semmle, 2023; Leong, 2022). However, these techniques face challenges in scalability and applicability. Fuzzing does not scale to large applications, and fuzzing parts of applications requires manually crafting fuzz drivers. Moreover, it is hard to use on large critical programs with complex inputs, such as network servers, GUI-based programs, embedded firmware, boot loaders, and system services. On the other hand, static analysis relies heavily on manual API specifications, and skillfully crafted heuristics to balance precision and scalability. In light of these challenges, GitHub pays a bounty of over 7K USD for each CodeQL query that can find new critical security bugs (GitHub, 2023).
Large Language Models (LLMs), including pre-trained models such as GPT-4 and CodeLlama, have made remarkable advances in code-related tasks in a relatively short period. Such tasks include code completion (Chen et al., 2021), automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024). These results clearly show the promise of LLMs, opening up a new direction for exploring advanced techniques. Hence, an intriguing question is whether the state-of-the-art pre-trained LLMs can also be used for detecting security vulnerabilities in code.
To develop LLM-based solutions, an important first step is to systematically evaluate the ability of LLMs in detecting known vulnerabilities. This is especially important in light of the rapidly evolving landscape of LLMs in three aspects: scale, diversity, and applicability. First, scaling these models to ever larger numbers of parameters has led to significant improvements over previous generations in their capabilities—a phenomenon termed as emergent behavior (Wei et al., 2022a). For instance, GPT-4, which is presumably orders of magnitude larger than its 175-billion predecessor GPT-3.5, significantly outperforms GPT-3.5 on a wide range of code-understanding tasks (Bubeck et al., 2023). Second, the diversity of LLMs has grown rapidly and now includes not only proprietary general-purpose ones such as GPT-4 but also open-sourced LLMs such as CodeLlama (Rozière et al., 2023) and StarCoder (Li et al., 2023) that are specialized for code. Finally, the reasoning capabilities of LLMs (and hence their applicability) may vary significantly across different programming languages. All these factors open up a large exploration space for applying LLMs to the challenging task of vulnerability detection.
Our Work. We conduct the first comprehensive study of using LLMs for detecting security vulnerabilities. We study five state-of-the-art LLMs, including proprietary models such as GPT-3.5 and GPT-4, and open-source models like CodeLlama. We evaluate these models on five popular security vulnerability datasets.
We design a set of three prompting strategies for LLMs to elicit increasingly sophisticated forms of reasoning and explanations. Our simplest prompting strategies include the Basic prompt, which simply asks an LLM to check for any vulnerabilities in the given code and the CWE specific prompt, which asks the LLM to check for a specific class of vulnerabilities or CWEs (such as Buffer Overflows).
A significant limitation of static vulnerability detection tools is the requirement of building the target project to enable them to find bugs. Further, they also require concrete specifications of APIs (e.g., sources, sanitizers, and sinks). In contrast, LLMs have an internal model of APIs already seen during the pre-training phase and hence they do not require compiled or complete codebases to run. Inspired by this insight, we additionally design a new prompting strategy, called Dataflow analysis-based prompt, that simulates a source-sink-sanitizer based dataflow analysis on the target code snippet before predicting if it is vulnerable. The Dataflow analysis-based prompt, similar to a classical dataflow-based static analyzer asks the LLM to first infer the sources, sinks, and sanitizers in the code snippet and check for any unsanitized data flows between the source and sink. This style of prompting is similar to chain-of-thought reasoning and simulates step-by-step reasoning by LLMs.
A key challenge in evaluating LLMs is data leakage: LLMs may perform well on a dataset because such samples were already present in their pre-training data. Hence, we implement three semantics-preserving adversarial attacks for code and evaluate whether they significantly impact the performance of LLMs.
Research Questions. We study the following research questions:
RQ1: How do different pre-trained LLMs perform in detecting security vulnerabilities across different languages and datasets? (Section §3.1)
RQ2: How do different prompting strategies affect the performance of LLMs? (Section 3.2)
RQ3: How does the performance of LLMs vary across different vulnerability classes? (Section 3.3)
RQ4: Can adversarial attacks impact the performance of LLMs in detecting vulnerabilities? (Section §3.4)
RQ5: How do LLMs compare to state-of-the-art static analysis tools? (Section 3.5)
RQ6: How do LLMs compare to state-of-the-art deep-learning-based tools? (Section 3.6)
Results and Findings. We choose five state-of-the-art LLMs: GPT-4, GPT-3.5, CodeLlama-34B, CodeLlama-13B, and CodeLlama-7B and evaluate their effectiveness in detecting security vulnerabilities across five vulnerability datasets: OWASP (OWASP Benchmark Suite, 2023), Juliet Java (Juliet Java, 2023), Juliet C/C++ (Juliet C/C++, 2023), CVEFixes Java (Bhandari et al., 2021), and CVEFixes C/C++ (Bhandari et al., 2021). These datasets cover 5000 vulnerable and non-vulnerable code samples, across 25 CWEs. Our findings are summarized as follows:
LLMs across all sizes have modest vulnerability detection ability, with an average accuracy of only about 60% across all datasets.
@@
Using prompts that focus on detecting specific CWEs improves the performance of LLMs. The dataflow analysis-based prompt further improves results for larger LLMs by allowing step-by-step reasoning. The improvement is significant for real-world datasets (an increase of up to 0.25 F1 score). However, we observe that LLMs often infer the relevant source, sink, and sanitizers correctly but fail in end-to-end logical reasoning.
LLMs are relatively better at detecting vulnerabilities that require local reasoning without complex code structures (such as Integer Overflow and NULL Pointer Dereference). However, they struggle to detect more complex vulnerabilities in real-world programs.
Larger LLMs like GPT-4 show mild degradation in performance when subject to adversarial attacks on code (average drop 8.6%). In contrast, smaller LLMs, like CodeLlama-7B, show a significant drop in performance (average drop 39%) in the presence of perturbations.
LLMs lag behind state-of-the-art static analysis tools like CodeQL in average accuracy (by 15%) and precision (by 26%) across datasets. However, LLMs like GPT-4 show some promising partial abilities, e.g., by identifying source/sink specifications or doing context-based reasoning using natural language information, which can potentially be useful for static analysis.
We obtain mixed results comparing LLMs to prior deep-learning-based tools: while the deep-learning-based tools only slightly outperform LLMs on synthetic datasets (by 0.07 in F1 score), all approaches struggle equally on real-world datasets often producing close to 50% accuracy only (same as a random baseline).
@@
Contributions. To summarize, we make the following contributions in this paper:
@@
Empirical Study: We conduct the first large comprehensive study on how state-of-the-art LLMs perform in detecting security vulnerabilities across five datasets and two programming languages (C/C++ and Java).
Prompting Strategies: We design three prompting strategies for LLMs, inspired by the recent advances in natural language processing and traditional program analysis techniques, that elicit different reasoning styles from LLMs and also provide human-readable explanations for their predictions.
Robustness of LLMs: We study how the performance of LLMs is impacted by semantics-preserving adversarial attacks on code.
@@
Comparison with other vulnerability detection tools: We contrast the performance of LLMs against popular static analysis and deep-learning-based vulnerability detection tools.
@@
2. Approach
2.1. Datasets
For our study, we select five diverse vulnerability datasets from two languages: C++ and Java. For each language, we also select both synthetic and real-world benchmarks. Table 1 presents the details of each dataset, such as the dataset size, programming language, number of vulnerable and non-vulnerable samples, and the number of unique CWEs. We describe each dataset next.
@@ 2.1. Datasets
| OWASP(OWASP Benchmark Suite, 2023) | Java | 2740 | 1415/1325 | 11 | |
| SARD Juliet (C/C++) (Juliet C/C++, 2023) | C/C++ | 81,280 | 40,640/40,640 | 118 | |
| SARD Juliet (Java) (Juliet Java, 2023) | Java | 35,940 | 17,970/17,970 | 112 | |
| CVEFixes (Bhandari et al., 2021) | C/C++ | 19,576 | 8223/11,347 | 131 | |
| CVEFixes (Bhandari et al., 2021) | Java | 3926 | 1461/2465 | 68 |
2.1.1. OWASP (Synthetic)
The Open Web Application Security Project (OWASP) benchmark (OWASP Benchmark Suite, 2023) is a Java test suite designed to evaluate the effectiveness of vulnerability detection tools. Each test represents a synthetically designed code snippet containing a security vulnerability. OWASP contains 2740 test cases representing 11 unique classes of security vulnerabilities (also known as Common Weakness Enumeration or CWE).
2.1.2. Juliet (Synthetic)
Juliet (Boland and Black, 2012) is a widely-used vulnerability dataset developed by NIST. Juliet comprises thousands of synthetically generated test cases representing various known vulnerability patterns. It contains 81,280 C/C++ programs covering 118 unique CWEs, and 35,940 Java programs covering 112 unique CWEs. For our paper, we use the latest version, Juliet 1.3 (Black and Black, 2018; Juliet C/C++, 2023; Juliet Java, 2023).
2.1.3. CVEFixes (Real-World)
Bhandari et al. (Bhandari et al., 2021) curated a dataset, known as CVEFixes, from 5365 Common Vulnerabilities and Exposures (CVE) records from the National Vulnerability Database (NVD). From each CVE, they automatically extracted the vulnerable and patched versions of each method in open-source projects, along with extensive meta-data such as the corresponding CWEs, project information, and commit data. CVEFixes consists of methods extracted from 5495 vulnerability-fixing commits. These methods span multiple programming languages such as C/C++, Java, Python, and JavaScript. For our work, we extracted all C/C++ and Java methods from CVEFixes. We collected 19,576 C/C++ and 3926 Java methods (both vulnerable and non-vulnerable), covering 131 and 68 different CWEs, respectively.
While many real-world datasets have been proposed in the literature, we needed a dataset for our study that 1) contains vulnerability metadata such as CVE and CWE IDs, 2) is two-sided, i.e., contains both vulnerable and non-vulnerable code samples, and 3) covers multiple languages such as Java and C/C++. Table 2 shows a comparison of existing real-world vulnerability datasets. We selected CVEFixes because it is the only dataset that fits our criteria.
| Dataset | Languages | CVE Metadata | Two-Sided | Multi-Lang |
| BigVul (Fan et al., 2020) | C/C++ | ✓ | ✗ | ✗ |
| Reveal (Chakraborty et al., 2021) | C/C++ | ✗ | ✗ | ✗ |
| DiverseVul (Chen et al., 2023) | C/C++ | ✗ | ✓ | ✗ |
| DeepVD (Wang et al., 2023) | C/C++ | ✗ | ✗ | ✗ |
| CVEFixes (Bhandari et al., 2021) | C/C++, Java, … | ✓ | ✓ | ✓ |
2.2. Metrics

### v3

@@ Abstract
Security vulnerabilities in modern software are prevalent and harmful. While automated vulnerability detection techniques have made promising progress, their scalability and applicability remain challenging. The remarkable performance of Large Language Models (LLMs), such as GPT-4 and CodeLlama, on code-related tasks has prompted recent works to explore if LLMs can be used to detect security vulnerabilities. In this paper, we perform a more comprehensive study by examining a larger and more diverse set of datasets, languages, and LLMs, and qualitatively evaluating detection performance across prompts and vulnerability classes. Concretely, we evaluate the effectiveness of 16 pre-trained LLMs on 5,000 code samples—1,000 randomly selected each from five diverse security datasets. These balanced datasets encompass synthetic and real-world projects in Java and C/C++ and cover 25 distinct vulnerability classes.
Our results show that LLMs across all scales and families show modest effectiveness in end-to-end reasoning about vulnerabilities, obtaining an average accuracy of 62.8% and F1 score of 0.71 across all datasets. LLMs are significantly better at detecting vulnerabilities that typically only need intra-procedural reasoning, such as OS Command Injection and NULL Pointer Dereference. Moreover, LLMs report higher accuracies on these vulnerabilities than popular static analysis tools, such as CodeQL.
@@
We find that advanced prompting strategies that involve step-by-step analysis significantly improve performance of LLMs on real-world datasets in terms of F1 score (by up to 0.18 on average). Interestingly, we observe that LLMs show promising abilities at performing parts of the analysis correctly, such as identifying vulnerability-related specifications (e.g., sources and sinks) and leveraging natural language information to understand code behavior (e.g., to check if code is sanitized). We believe our insights can motivate future work on LLM-augmented vulnerability detection systems.
I Introduction
@@
Security vulnerabilities afflict software despite decades of advances in programming languages, program analysis tools, and software engineering practices. Even well-tested and critical software such as OpenSSL, a widely used library for applications that provide secure communications, contains trivial buffer overflow vulnerabilities, e.g., [1] and [2]. A recent study by Microsoft showed that more than 70% of vulnerabilities are still caused by well-understood memory safety issues [3]. This is alarming given the rapidly growing size and complexity of modern software systems, encompassing numerous programs, libraries, and modules that interact with each other. Hence, we need major technical advances to effectively detect security vulnerabilities in such complex software.
@@
Traditional techniques for automated vulnerability detection, such as fuzzers [4], and static analyzers such as CodeQL [5] and Semgrep [6] have made promising strides. For example, in the last two years, researchers found over 300 security vulnerabilities through custom CodeQL queries [7, 8]. However, these techniques face challenges in scalability and applicability. Fuzzing requires manually crafted fuzz drivers and does not scale to large critical programs with complex inputs, such as network servers, embedded firmware, and system services. On the other hand, static analysis relies heavily on manual API specifications, and skillfully crafted heuristics to balance precision and scalability. Until recently, GitHub paid a bounty of over 7K USD for each CodeQL query that found new critical security bugs [9].
| Research Questions | Findings | |
| RQ1: How do different pre-trained LLMs perform in detecting security vulnerabilities across different languages and datasets? (Section §III-A) | ✓ LLMs across all sizes report a mean accuracy of about 62.8% and a mean F1 score of 0.71 across all datasets. ✗ Average accuracy on real-world datasets is 10.5% lower than that on synthetic datasets. ✓ In stark contrast to other domains, smaller models such as Qwen-2.5-14B and Qwen-2.5-32B report higher accuracies on the real-world datasets than much larger models such as GPT-4. | |
@@
| RQ2: How do different prompting strategies affect the performance of LLMs? (Section III-B) | ✓ Using prompts that focus on detecting specific CWEs improves the performance of LLMs. ✓ The Dataflow analysis-based prompt prompt further improves results for larger LLMs with an increase of up to 0.18 in F1 score on real-world datasets. ✓ We also observe that LLMs often infer the correct sources/sinks/sanitizers but fail in end-to-end reasoning. | |
| RQ3: How does the performance of LLMs vary across different vulnerability classes? (Section III-C) | ✓ LLMs are better at detecting local vulnerabilities that require no global context across datasets (OS Command Injection, NULL Pointer Dereference, Out-of-bounds Read/Write, etc.). ✗ LLMs struggle to detect vulnerabilities that require additional context or reasoning about complex data structures (Out-of-bounds Read/Write with C++ structs and pointers). ✓ Certain LLMs are very good at detecting specific CWEs across datasets (Llama-3.1-70B on OS Command Injection, DeepSeekCoder-7B on NULL Pointer Dereference, etc.). | |
| RQ4: How do LLMs compare to state-of-the-art static analysis tools? (Section III-D) | ✗ LLMs report lower overall accuracies than CodeQL on all synthetic datasets. ✓ LLMs report higher accuracies than CodeQL on certain vulnerability classes across datasets (Path Traversal, OS Command Injection, etc). CodeQL reports higher accuracies on Integer Overflow across datasets. | |
| RQ5: How do LLMs compare to state-of-the-art deep-learning-based tools? (Section III-E) | ✓ Deep Learning(DL)-based tools such as DeepDFA [10] and LineVul [11] report accuracies similar to Qwen-2.5-32B on CVEFixes C/C++ even after being trained on in-distribution samples whereas Qwen-2.5-32B reports higher F1 scores. ✓ DL-based tools report lower accuracies and F1 scores than LLMs when trained and evaluated on different datasets. ✓ LLMs provide natural language explanations for their predictions while DL-based tools provide binary scores and line numbers that are often difficult to interpret. |
Large Language Models (LLMs), including pre-trained models such as GPT-4 and CodeLlama, have made remarkable advances in code-related tasks in a relatively short period. Such tasks include code completion [12], automated program repair [13, 14, 15], test generation [16, 17], code evolution [18], and fault localization [19]. These results clearly show the promise of LLMs, opening up a new direction for exploring advanced techniques. Hence, an intriguing question is whether the state-of-the-art pre-trained LLMs can also be used for detecting security vulnerabilities in code.
To develop LLM-based solutions, an important first step is to systematically evaluate the ability of LLMs in detecting known vulnerabilities. This is especially important in light of the rapidly evolving landscape of LLMs in three aspects: scale, diversity, and applicability. First, scaling these models to ever larger numbers of parameters has led to significant improvements in their capabilities[20]. For instance, GPT-4, which is presumably orders of magnitude larger than its 175-billion predecessor GPT-3.5, significantly outperforms GPT-3.5 on a wide range of code-understanding tasks [21]. Second, the diversity of LLMs has grown rapidly and now includes not only proprietary general-purpose ones such as GPT-4 but also open-source code-specific LLMs such as CodeLlama [22] and StarCoder [23]. Finally, the reasoning capabilities of LLMs (and hence their applicability) may vary significantly across different prompting strategies and programming languages. All these factors open up a large exploration space for applying LLMs to the challenging task of vulnerability detection.
| Study Features | [24] | [25] | [26] | [27] | [28] | Our Work |
| Languages | C/C++ | C/C++ | C/C++ | C,Py | C/C++ | C/C++,Java |
| #Samples | 368 | 347 | 100 | 96 | 25.9 | 5000 |
| #CWEs | 25 | N/A | N/A | 8 | 140 | 25 |
| #LLMs (1B parameters) | 3 | 16 | 11 | 5 | 4 | 16 |
| Comparison of various LLMs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| Qualitative prompt analysis | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ |
| CWE-wise analysis | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ |
| Comparison with SA tools | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ |
| Comparison with DL tools | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ |
Our Work. We study the vulnerability detection capabilities of 16 state-of-the-art LLMs across different scales and families, including proprietary models such as Gemini and GPT-4, and open-source models like CodeLlama and Qwen. We evaluate these models on five popular security vulnerability datasets across two languages, C/C++ and Java, and 25 vulnerability classes.
We first study how LLMs perform on the task of vulnerability detection using three prompting strategies and how these strategies qualitatively compare against each other. We also attempt to identify vulnerability classes that most benefit from the use of LLMs and these prompting techniques. Our simplest prompting strategies include the Basic prompt, which simply asks an LLM to check for any vulnerabilities in the given code and the CWE specific prompt, which asks the LLM to check for a specific class of vulnerabilities or CWEs (such as Buffer Overflows). Inspired by the success of static analysis tools like CodeQL that use dataflow analysis to detect vulnerabilities, we design a prompting strategy called Dataflow analysis-based prompt. This prompt asks the LLM to simulate a source-sink-sanitizer based dataflow analysis on the target code snippet before predicting if it is vulnerable.
We next compare LLMs with existing vulnerability detection tools, namely static analysis-based CodeQL and two deep learning-based techniques, DeepDFA [10] and LineVul [11]. As discussed earlier, static vulnerability detection tools are limited by the need for concrete API specifications and require compiling / building entire target projects before detection. Pre-LLM deep learning-based approaches such as DeepDFA [10] and LineVul [11] attempt to mitigate some of these limitations through fine-tuned neural representations of code. On the other hand, LLMs do not require the compilation of complete projects as they can be prompted to analyze partial code snippets. Moreover, they already have an internal model of APIs through pre-training and do not need to be trained on large datasets from scratch. We analyze the benefits and shortcomings of LLMs over CodeQL, including vulnerability classes where they outperform each other. We also study how they compare against the deep learning based-approaches in terms of generalization across datasets.
Comparison with other studies. There are other studies that also evaluate the effectiveness of LLMs on the task of vulnerability detection [24, 25, 26, 27, 28]. Table II presents a comparison of our work with these studies. We present our study as the most comprehensive on this topic for the following reasons:
Size and diversity of the datasets: We curate a dataset of 5K samples, with equal number of vulnerable and non-vulnerable snippets. The only larger dataset is from [28] but only 695 of their 25.9K samples are vulnerable. Furthermore, our study is the first to include Java code.
Comparison with non-LLM-based tools: Our study is the first to compare LLMs with CodeQL and specialized Deep Learning-based tools. Moreover, we also find vulnerability classes where LLMs outperform CodeQL and vice versa which is useful for deployment.
Qualitative analysis of prompts and vulnerability classes: While other studies quantitatively compare prompting strategies, we also attempt to qualitatively identify the benefits of various prompt elements. Furthermore, we identify partial capabilities offered by some of these prompts that can be leveraged in LLM-based detection tools. We also identify vulnerability classes where LLMs perform well.
@@
Contributions. Our research questions and key findings are summarized in Table I. To summarize, we make the following contributions in this paper:
@@
Empirical Study: We conduct the largest comprehensive study on how state-of-the-art LLMs perform in detecting security vulnerabilities across 5000 samples from five datasets, two programming languages (C/C++ and Java) and covering 25 unique vulnerability classes.
@@
Comparison with other vulnerability detection tools: We contrast the performance of LLMs against popular static analysis and deep-learning-based vulnerability detection tools. We also identify vulnerability classes where LLMs perform better/worse than some of these tools.
Qualitative comparison of prompting strategies: We quantitatively and qualitatively compare three prompting strategies, including a novel prompt inspired by dataflow analysis-based vulnerability detection.
@@
II Approach
II-A Datasets
For our study, we select five diverse synthetic/real-world vulnerability datasets from two languages: C++ and Java. Table III presents the details of each dataset, such as the dataset size, programming language, number of vulnerable and non-vulnerable samples, and the number of unique CWEs. The synthetic benchmarks, OWASP and Juliet, allow for easy comparison with CodeQL and the real-world benchmarks, CVEFixes, are useful for evaluating practical utility. While many real-world datasets have been proposed in the literature, we selected CVEFixes because it is the only dataset that 1) contains vulnerability metadata such as CVE and CWE IDs, 2) is two-sided, i.e., contains both vulnerable and non-vulnerable code samples, and 3) covers multiple languages such as Java and C/C++. Table IV shows a comparison of existing real-world vulnerability datasets. We briefly describe each of the selected datasets next:
@@ II-A Datasets
| OWASP[29] | Java | 2740 | 1415/1325 | 11 | |
| SARD Juliet (C/C++) [30] | C/C++ | 81,280 | 40,640/40,640 | 118 | |
| SARD Juliet (Java) [31] | Java | 35,940 | 17,970/17,970 | 112 | |
| CVEFixes [32] | C/C++ | 19,576 | 8223/11,347 | 131 | |
| CVEFixes [32] | Java | 3926 | 1461/2465 | 68 |
II-A1 OWASP (Synthetic)
The Open Web Application Security Project (OWASP) benchmark [29] is a Java test suite designed to evaluate the effectiveness of vulnerability detection tools. Each test represents a synthetically designed code snippet containing a security vulnerability.
II-A2 Juliet (Synthetic)
Juliet [33] is a widely-used vulnerability dataset developed by NIST with thousands of synthetically generated test cases from various known vulnerability patterns.
II-A3 CVEFixes (Real-World)
Bhandari et al. [32] curated a dataset, known as CVEFixes, from 5365 Common Vulnerabilities and Exposures (CVE) records from the National Vulnerability Database (NVD). From each CVE, they automatically extracted the vulnerable and patched versions of each method in open-source projects, along with extensive meta-data such as the corresponding CWEs, project information, and commit data. These methods span multiple programming languages but we only consider the C/C++ and Java methods in this work.
II-B Metrics